OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the open sentence 4 – 2|n + 6| ≥ 2? A. –5 < n or n < –7 B. n + 6 ≤ 1 and n + 6 ≥ –1 C. n + 6 ≤ –1 and n + 6 ≥ 1 D. n > –5 or n < –7

OpenStudy (jdoe0001):

\(\bf 4 - 2|n + 6| \ge 2\implies -2|n + 6| \ge 2+4\implies -2|n + 6| \ge 6\\ \quad \\ \textit{recall, in an inequality multiplying/dividing/exponentializing}\\ \textit{ by a negative number, we have to }flip \textit{ the inequality sign}\qquad thus\\ \quad \\ \cfrac{-2|n + 6|}{-2} \ge \cfrac{6}{-2}\implies |n+6|\le -3\\ \quad \\ |n+6|\le -3\implies \begin{cases} +(n+6)\le -3\\ \quad \\ \bf -(n+6)\le -3 \end{cases}\) solve those 2 scenarios, what do they give you? again, do not forget to "flip the inequality" sign when multiplying or dividing or exponentializing by a NEGATIVE number

OpenStudy (anonymous):

wow thats amazing thank you so much that helps so much thank you

OpenStudy (jdoe0001):

yw